ABC Company has set the following standards in order to produce one unit of its single product: standard quantity standard price direct materials 2.7 yards ?? per yard direct labor 2 hours $14 per hour variable overhead 2 hours ?? per hour During August, ABC Company spent $133,770 to purchase direct materials and had direct labor totaling $189,440. During August, ABC Company used 16,000 yards of direct materials in the production of units. ABC had 2,500 yards of materials in its August 1 direct material inventory and had 1,200 yards yards remaining in its direct materials inventory at August 31. ABC Company reported the following variances for August: Direct material price variance .............. $12,495 unfavorable Direct labor rate variance .................. $24,240 unfavorable Total direct labor variance ................. $6,560 favorable Variable overhead spending variance ......... $1,900 unfavorable Variable overhead efficiency variance ....... $20,900 favorable Calculate ABC Company's actual variable overhead cost for August.
Answer #1

Solution:

Actual direct labor cost = $189,440

Standard direct labor cost = Actual direct labor cost + Total favorable direct labor variance = $189,440 + $6,560 = $196,000

Nos of units produced = $196,000 / $28 = 7000 units

Standard hours for actual production = 7000*2 = 14000 hours

Direct labor efficiency variance = $24,240 + $6,560 = $30,800 Favorable

(SH - AH) *SR = $30,800

(14000 - AH) * $14 = $30,800

Actual labor hours = 11800 hours

Variable overhead efficiency variance = $20,900 F

(SH - AH) * SR = $20,900

(14000 - 11800) * SR = 20,900

Standard variable overhead rate = $9.50 per hour

Standard cost of variable overhead = 14000 * $9.50 = $133,000

Actual variable overhead cost = Standard variable overhead cost + Unfavorable variable overhead spending variance - Favorable variable overhead efficiency variance

= $133,000 + $1,900 - $20,900

= $114,000